Point charges q1=50μC and q2=−25μC are placed 1.0 m apart. What is the force on a third charge q3=20μC. Where must q3 of the preceding problem be placed so that the net force on it is zero?
Answer #1

let us consider that point is to the right of q2 at a distance d from it. at that point electric field due to q1 must be equal to electric field due to to q2

k q1 / ( 1 + d)^2 = k q2 / d^2

sqrt (q1 /q2) = 1/d + 1

d = 2.41 cm to the right of q2

=======

Comment in case any doubt, will reply for sure.. Goodluck